A gas has a solubility in water of 11.9 g/L at 15°C and 505 kPa of pressure. What is the pressure of the gas, in atm., (at the same temperature) if the solubility of the gas is 20 mg/L?

Your answer


734 grams of lithium sulfate are dissolved to make 2500.0 mL of solution. What is the Molarity of this Solution

Your answer


Describe the process you would follow to prepare 120 ml of a 2.0 M solution of HCl if you started with a 6.0 M stock solution.

Your answer


What is the percent (mass/volume) if 12 g of KCl is added to 50 mL of water to make 93 mL of solution?

Your answer


If 75 g. of Potassium Chloride (ionic compound) is dissolved in 250 grams of water, what will be the freezing point of the solution? Kf = 1.86

Your answer


The vapor pressure of pure water is 3.17 kPa. What will be the vapor pressure of a solution containing 13 g. of potassium permanganate (KMnO4) and 750 g of water?


1
Expert's answer
2021-04-05T07:51:37-0400

1. P1= 505Kpa= 505000Pa= 4.98atm

m1= 11.9g/L

T1=288K

P2=?

m2=20mg/L

From ideal gas equation

PV= nRT

PV= m/M x RT

We can see that P is directly proportional to mass concentration, so we can set the equation

P1/m1=P2/m2

P2=P1m2/m1=4.98x20/11.9=8.37g/L


2. Molar mass of Li2SO4=110g/mol

V=2500ml=2.5L

m=734g

Mole= mass/molar mass=734/110=6.67mol

Molarity= moles/litres= 6.67/2.5=2.67M


3. C1= 6.0M

V1= ?

C2= 2.0M

V2= 120ml

C1V1=C2V2

V1= C2V2/C1= 2.0x120/6.0= 40ml

To prepare this dilute solution, pipette 40ml of the concentrated HCl into a standard flask and make it up to 1Liter.


4. %w/V= 12g/50ml x 100%=24%w/V


5. Mass of KCl= 75g

Moles of KCl= mass/molar mass= 75/74.5=1.007mol

Mass of water= 250g= 0.25Kg

Molality= moles of solute/Kg of solvent

Molality= 1.007/0.25= 4.03M

It follows that

∆T= MKf

Where ∆T is the boiling point

∆T= 4.03x1.86=7.49°C


6. Vapour pressure of pure water Pi= 3.17Kpa

Mass of water= 750g

Mole of water= 750/18= 41.7mol

Mass of KMnO4= 13g

Molar mass of KMNO4= 158g/mol

Mole of KMnO4= 13/158= 0.082mol

Mole fraction of water, Xi= 41.7/(41.7+0.082) = 0.998

Pressure of solution, Ps= ?

From Raoult's law

Pi=XiPs

Ps= Pi/Xi= 3.17/0.998= 3.176KPa